carmelle:

Write the equation of a line in standard form that has x-intercept (−P, 0) and y-intercept (0, −R). A. Px − Ry = −PR B. Px − Ry = PR C. Rx + Py = −PR D. Rx + Py = PR

carmelle:

i don't understand...

Joe348:

\[\frac{ y2-y1 }{ x2-x1}\] you understand you are using this equation right?

carmelle:

yea, i already put it in y = mx + b format and got: slope:\[-\frac{ R}{ P }\] \[y = - \frac{ R }{ P }x~-R\]

Joe348:

Okay so multiply the equation by P ;-;

carmelle:

im lost...

Joe348:

You have to re-arrange the equation so to do that you multiply it by P, you see the p on the equation \[-\frac{ r }{p}x\] multiply everything by P on each side

Joe348:

once you do that, it'll be \[py=-rx+rp\]

Joe348:

But your not totally finished, as you can see that's not of the answer choices, so you need to reword it

Joe348:

@joe348 wrote:
once you do that, it'll be \[py=-rx+rp\]
Are you able to finish it yourself?

carmelle:

\[rx ~+~ py =~pr\] i got D?

Joe348:

Yeah, nice

carmelle:

tysm :)

Joe348:

np ;) glad to help now help me lol

carmelle:

D ended up being wrong :')

Joe348:

oh im so sorry

carmelle:

nah, its fine, dw